I have a right triangle with side a=7 and side b=5. What is the value of sin A?
Answer by stanbon(48564)

I have a right triangle with side a=7 and side b=5. What is the value of sin A?
------------------
You need to solve for "c" which I assume is the hypotenuse.
c = sqrt(7^2 + 5^2)
c = sqrt(49 + 25)
c = sqrt(74)
-------
If c is the hypotenuse, sin(A) = a/c = 7/sqrt(74)
========================================================
Cheers,
Stan H.